OpenStudy (anonymous):

(-2x^2+6x +1)-2(4x^2-3x+1)=

OpenStudy (anonymous):

(-2x^2+6x + 1)-2(4x^2-3x +1) =(-2x^2+6x +1)-8x^2+6x -2 =-2x^2 +6x + 1-8x^2 +6x -2 =-2x^2-8x^2+6x+6x +1-2 =-10x^2 +12x -1